OpenStudy (madhatter):

so in a math equation that looks for example like this 2b+5q=9+4q does both q's= the same thing

OpenStudy (phi):

yes. you have 5 q's on one side and 4 q's on the other side of the equation they are "similar terms"

OpenStudy (madhatter):

so 5 xq and 4x q =? but if i multplid 4 xq and 5x q = is the same what ever nuber it ends up to be?

OpenStudy (phi):

if you have an equation with one unknown, for example: 20+5q=9+4q we can combine like terms: add -4q to both sides 20 + 5q - 4q = 9 +4q - 4q on the right we have 4 q's take away 4 q's leaving 0 q's or just 0 20 + 5q - 4q = 9 +0 9+0 is just 9, so this is 20 + 5q - 4q = 9 on the left side you have 5 q's take away 4 q's leaving 1 q 20 +1q =9 (1q is normally written as "q") 20 + q = 9 now add -20 to both sides and you get q= 9-20 q= -11 however, if you have 1 equation and 2 unknowns , like your equation we cannot solve q for a specific number. But we can simplify things 2b+5q=9+4q q= 9 - 2b (do you see how to do that ?)